OpenStudy (turingtest):

Definite integral\[\int_0^1\frac x{x^3+1}dx\]

OpenStudy (turingtest):

I've had this one on my head for a couple days now, so whoever can help me gets mad respect :)

OpenStudy (experimentx):

you can do that using partial fraction. but this is not quite interesting.

OpenStudy (turingtest):

the solution from PF I got seemed drastically different than wolf's elegant answer, though I could have just been unable to manipulate the algebra.

OpenStudy (turingtest):

Is PF really the only way I guess is my real question

OpenStudy (turingtest):

it seems like there should be a trick

OpenStudy (experimentx):

yeah ... very short and elegant method.

OpenStudy (turingtest):

I know, that is why I want something besides hacking away with PF which gives http://www.wolframalpha.com/input/?i=integral+x%2F%281%2Bx%5E3%29

OpenStudy (turingtest):

okay now I see they are the same, so either that was a dumb question, or there is a trick I want to know

OpenStudy (experimentx):

there might be.

OpenStudy (experimentx):

trig subs might work. but this seems more nasty.

OpenStudy (dape):

Yeah, PF would definitely work. Probably won't be able to do much, but I'll try to poke around and see if I can find a nicer method.

OpenStudy (dape):

Muhahaha

OpenStudy (amistre64):

... id have to brute away at it :/

OpenStudy (dape):

It turns out after researching the integrand for a bit that it has an \(extremely\) nice McLaurin series. I don't have the energy to type out the entire derivation, but you can check it's: \[ \frac{x}{x^3+1}=x-x^4+x^7-x^{10}+...=x\sum_{n=0}^\infty(-1)^nx^{3n} \] Throwing in this in the integral and doing a partial integration we get: \[ \int_0^1{\frac{x}{x^3+1}\, dx}=\int_0^1{x\sum_{n=0}^\infty(-1)^nx^{3n}\,dx} \\ =\sum_{n=0}^\infty \frac{(-1)^n}{3n+1}-\sum_{n=0}^\infty(-1)^n\int_0^1{x^{3n}\,dx} \\ =\sum_{n=0}^\infty\frac{(-1)^n}{3n+1}\left(1-\frac{1}{3n+2}\right) \\ =\sum_{n=0}^\infty\frac{(-1)^n}{3n+2} \] Now that sum is a bit harder to calculate, it involves zeta-functions in combination with some subtitutions, but at least it comes out the same as wolf's answer :p

OpenStudy (dape):

I think it's a more elegant solution, despite it involving some pretty high level stuff.

OpenStudy (dape):

That was fun, when I first saw that MacLaurin series I got some serious chills down my spine.

OpenStudy (dape):

Oops, typo, the integral term on second line should be: \[ -\sum_{n=0}^\infty{\frac{(-1)^n}{3n+1}}\int_0^1{x^{3n+1}\,dx} \]

OpenStudy (anonymous):

heh, was gonna say, you lost an x!

OpenStudy (anonymous):

then i saw you just forgot it

OpenStudy (experimentx):

this seems pretty straight forward \[ \int_0^1{x\sum_{n=0}^\infty(-1)^nx^{3n}\,dx} = \sum_{n=0}^\infty(-1)^n \int_0^1x^{3n+1}\,dx = \sum_{n=0}^\infty {(-1)^n \over 3n+2}\] I get chills when i see these.

OpenStudy (anonymous):

there's some extra stuff in there...

OpenStudy (experimentx):

I'm not quite well with zeta functions ... looks impossible from Fourier series.

OpenStudy (anonymous):

whatever, if it's actually supposed to be there, I can't figure out why..

OpenStudy (dape):

Maybe I forgot something else, did it all on paper and copied it over.

OpenStudy (anonymous):

seems it's just as experimentx did it, what's with the first series and the "(1-"

OpenStudy (experimentx):

still ... i never though of this technique. I haven't been using this technique lately. I'll try to think of this way too in the future. gotta sleep ... nearly 3.30 am here.

OpenStudy (anonymous):

If you've gone McLaurin - then why not go full complex contour integration. The poles are obvious -1, +third-root(-1), -third-root(-1)

OpenStudy (dape):

Arent the poles \(-1,\sqrt[3]{-1},-(\sqrt[3]{-1})^2\)?

OpenStudy (anonymous):

Isn't this thing just a trig substitution?

OpenStudy (anonymous):

Never mind I see the x cubed. I can't see how to do this without partial fractions :/ .

OpenStudy (turingtest):

nice @dape that was definitely more what I was looking for, now this integral is interesting again :)

OpenStudy (anonymous):

Yes @dape these are the poles. Yet I am not certain there is an effective choice of contour so that we can find the needed contribution

OpenStudy (turingtest):

Contour integral are something I certainly learn, but I would then like to see it that way too.

OpenStudy (turingtest):

certainly need to*

OpenStudy (anonymous):